OpenStudy (anonymous):

A number is divided by 3. If 8 is subtracted from the quotient you get -6.

jimthompson5910 (jim_thompson5910):

"A number is divided by 3" means that we get x/3 Add on the fact that "If 8 is subtracted from the quotient you get -6", which means x/3 - 8 = -6 Solve this equation for x to find this unknown number.

OpenStudy (anonymous):

nopee your wrong ialready got the answer but thanks.

OpenStudy (anonymous):

x=2/3

jimthompson5910 (jim_thompson5910):

Well obviously I wasn't finished. That's just one of the steps to get to the answer.

OpenStudy (anonymous):

@jim_thompson5910's answer probably would have been, x=6.\[\frac{6}{3}-8=-6 \]